To be considered a trader in securities for tax purposes

43. To be considered a trader in securities for tax purposes, the trading activity must meet three general requirements f the following is not one of the three general requirements? 

  • Activity must be substantial. 

  • Transactions must be for a large amount. 

  • Transactions must be regular and continuous. 

  • Transactions must be intended for short-term swings 


44. Alex purchased 200 shares of ABC stock on June 2, 2016, for $1,200. In November 2017, ABC issued a 10% stock dividend. How many shares does Alex own now, and what is his basis? 

  • 220 shares, $1,080 basis. 

  • 220 shares, $1.200 basis. 

  • 220 shares, $1.320 basis. 

  • 220 shares, $1,452 basis 


45. Which of the following statements regarding the reporting of sales of stock is true? 

  • Short-term and long-term capital gains and losses are reported on Form 8949 and carried to Schedule D

  • Total capital gain is computed by combining net short-term gains and losses and net long-term gains and losses. 

  • Under certain circumstances, short-term and long-term capital gains and losses may be aggregated and reported directly on Schedule D. 

  • All of these statements are true.
Answer #1

Answer 43:

Correct answer is:

Transaction must be for a large amount.

Explanation:

'Transaction must be for a large amount' is not one of the three general conditions to be complied to be considered a trader in securities for tax purpose. (Source: https://www.irs.gov/taxtopics/tc429)

As such option B is incorrect. Other 3 options are incorrect since these 3 options are general conditions to be complied to be considered a trader in securities for tax purpose.

Answer 44:

Correct answer is:

220 shares, $1,200 basis

Explanation:

Number of Shares purchased on June 2, 2016 = 200

10% stock dividend issued in November = 200 * 10% = 20 shares

Total shares Alex owns = 200 + 20 = 220 shares.

His tax basis is purchase cost of 200 shares = $1,200

As such option B is correct and other options A, C and D are incorrect.

Answer 45:

Correct answer is:

All of these statements are true.

Explanation:

Total capital gain is computed by combining net short gains and losses and net long-term gains and losses.

Under certain circumstances short and long term capital gains and losses may be aggregated and reported directly on schedule D.

Since 2011, IRS introduced Form 8949 that some tax payers have to fill along with schedule D to report short term and long term capital gains and losses.

As such options A, B and C are all correct. But option D is most appropriate.